# A Systematic Literature Review of Software Engineering Research on Jupyter Notebook

### v1

@@ 5.1.3 Reproducibility
Reproducibility ensures that other notebook users can get the same results from a notebook by following the same steps, using the same data, and working in the same environment [3]. Reproducibility differs from code reuse and cloning, which typically involves adapting code or directly copying code into different environments and applying the code to different data to meet different objectives of users, resulting in outputs that are different from the original notebook.
@@
Reproducibility is known as one of the key promises of Jupyter notebooks, as they are often used to share results with others, allowing users to trace the steps from raw data to final results [72]. However, achieving reproducibility remains challenging due to issues such as out-of-order execution and repetitive execution of the same cell [177]. These issues differ from traditional programming, where code execution is typically linear (from top to bottom), ensuring a clear and consistent execution flow. For example, notebook users might execute the same cell multiple times and only save the latest execution state, causing cells located at the notebook’s beginning to be executed after later cells. This nonsequential execution can confuse others trying to reproduce results, affecting notebook reproducibility [177, 178].
@@
A study by Pimentel et al. [115] analyzed more than 1.4 million Jupyter notebooks from GitHub and found significant reproducibility issues. The study noticed that only 25% of the valid notebooks executed without errors under straightforward conditions, and only 5% reproduced results identical to their original outputs. Furthermore, Schröder et al. [143] revealed that only 14% of the available notebooks in published academic research on PubMed, [^11] https://pubmed.ncbi.nlm.nih.gov/ where Jupyter notebooks are used as implementation, were reproduced successfully. The study noticed the need for comprehensive documentation and containerization of notebooks, which supports the good practices presented by Rule et al. [133]. We have analyzed the relevant studies and presented a list of good and bad practices that affect the reproducibility of a notebook in Table 5.
@@
Researchers have demonstrated that the reproducibility of Jupyter notebooks can be significantly improved by resolving cell dependencies within the notebook’s code [177, 180]. Cell dependencies represent the relationships between variables, functions, and data in different code cells. Wang et al. [178] proposed a method to build a cell dependency graph using static analysis to accurately model these relationships and found that 83% of executable notebooks can be reproduced using their approach. Samuel and König-Ries [138] developed a visualization tool ReproduceMeGit to check the reproducibility of Jupyter notebooks from GitHub by extending the work from Pimentel et al. [115]. This tool can automatically install dependencies from requirements.txt or setup.py files and execute the notebook to provide highlights of the reproducibility study through a user interface.
@@ 5.1.4 Provenance
Provenance in notebooks refers to the detailed history of actions, executions, and dependencies that led to the current state of a notebook. It includes the sequence of cell executions, the modifications made to cells, variable dependencies, data flow, and intermediate outputs [68]. Unlike reproducibility which emphasizes achieving identical results under a consistent initial environment, provenance addresses the historical and iterative details present in the notebook’s creation and evolution. Managing provenance in Jupyter Notebook differs from traditional systems due to its iterative and non-linear nature. In a notebook, users can execute cells in any order, creating challenges in tracking the execution history and dependencies [69]. Additionally, Jupyter notebooks merge code, data, and outputs into a single file, whereas traditional systems follow a sequential execution process that generates output externally. As traditional systems offer clear execution pipelines and distinct output demonstrations, provenance management becomes more straightforward. In contrast, Jupyter notebooks require specialized tools for managing provenance to capture dynamic state changes and overwritten variables [68, 74].
@@
Researchers developed provenance detection techniques specialized for Jupyter Notebook to reuse notebooks by obtaining the original results. In the early stage of Jupyter Notebook usage, Pimentel et al. [117] proposed a notebook tool to overcome the limitations in provenance support by automatically capturing and analyzing the execution history and environment settings of code within notebooks. The study extends the noWorkFlow tool [102] developed for Python scripts to capture the provenance of scripts, including control flow information and library dependencies. The proof can be inferred by statically analyzing code cells on execution counts and cell positions within the notebooks [74]. A scheduler that adjusts the order of cell execution based on data dependencies detected during runtime can be used to improve the accuracy of provenance of Python code in notebooks [28].
@@
In addition to focusing on tools and execution kernels, researchers have developed extensions to capture provenance. MLProvLab [68] is a JupyterLab extension to track, manage, compare and visualize the provenance of computational experiments. It can capture the provenance at run-time by detecting dependency graphs. Another notebook extension, ProvBook, was designed to capture and visualize the provenance of notebook executions [137]. This extension automatically stores provenance information within the notebook’s metadata, detailing execution times, inputs, and outputs for each cell. It helps users to see how data and analysis have evolved over time and compare results, which helps to enable notebook reproducibility. Kery and Myers [66] developed an extension named Verdant that records the history of all previous changes in a notebook and quickly retrieves versions of a specific artifact from the existing versions of the entire document. It can help to compare multiple versions of different notebook artifacts, including code cells, tables, and images, which are absent in traditional version control systems like Git.
@@
Expanding on these efforts, Studtmann et al. [154] introduced Histree for automatic versioning and visual branching of notebook modifications. It organizes experiment histories into tree-based structures for easier navigation. Further extending provenance into interactive visual analytics, Gadhave et al. [37] proposed Persist, which captures interaction provenance across code and visualizations. Empirical evaluations showed improvements in analysis efficiency, accuracy, and reproducibility. Finally, Retrograde [54] integrates provenance tracking with fairness auditing. It provides real-time, context-aware notifications triggered by notebook events and maintains a dynamic data ancestry graph, helping users detect and address fairness and bias concerns during data preprocessing and modeling.
@@
Summary of code reuse and provenance Code reuse in Jupyter notebooks is essential for improving productivity and collaboration while addressing the challenges associated with their unique structure and iterative workflows. Users can more effectively leverage existing resources and maintain reusable notebooks through techniques such as code cloning, context-aware code search, provenance management, reproducibility practices, and advanced versioning tools. The development of innovative solutions, such as context-sensitive code search tools [84, 85], code dependency resolution systems [178], and provenance tracking extensions [68, 117, 54], highlights ongoing efforts to support effective code reuse and provenance. By integrating these solutions and best practices (see Table 5), notebooks can become more reusable and maintainable, thus establishing a stronger foundation for data science workflows and collaborative research.
@@ 5.2 Managing Computational Environment and Workflow
Managing the computational environment and workflow in Jupyter notebooks is crucial to ensure the successful execution and reproducibility of data science tasks [118, 92]. Unlike traditional programming environments, which are typically well defined and managed by integrated development environments or build systems, a Jupyter Notebook does not have a separate file to manage the configuration [72]. Traditional programming environments commonly include a configuration file (e.g., “requirements.txt” in Python) to ensure a consistent and reproducible setup for code execution. In contrast, Jupyter notebooks execute code cells independently, and each notebook has its own dependencies [25, 124]. This section explores various aspects of computational environment management, including empirical studies on notebook workflows, the impact of different environments on code behaviour, and the challenges associated with managing library dependencies.
@@ 5.2.1 Empirical Studies on Workflows
Lau et al. [78] analyzed 60 notebooks and summarized four main stages of a data science workflow: importing data into notebooks, writing and editing code, running the code to generate output, and publishing the results. However, workflows can vary. For example, notebooks often start with exploration, where users write lots of code to find interesting patterns in the data, and end with a presentation or explanation [124]. In addition, recent empirical analyses have shed further light on workflow diversity. Golendukhina and Felderer [43] studied 138,376 Kaggle notebooks and identified significant variations in data preprocessing practices correlated with user expertise, highlighting a notable gap between model-centric activities and actual data cleaning efforts. Similarly, Zou et al. [205] uncovered that data scientists frequently and manually experiment with alternative ML pipeline configurations (e.g., data preparation, model selection), emphasizing the limitations of managing and systematically exploring pipeline variations in traditional notebook environments. On the other hand, Psallidas et al. [118] identified two types of pipelines in Jupyter notebooks that are key components of data science workflows: explicit and implicit. Explicit pipelines utilize tools like sklearn.pipeline to define structured steps for tasks such as data pre-processing and model training. In contrast, implicit pipelines rely on ad hoc function calls with libraries such as Pandas to clean, merge, and visualize data [118]. Researchers have also suggested ways to map workflows to better understand how notebooks work. One approach uses directed acyclic graphs to show the flow of data and tasks within a notebook [130]. These graphs can highlight where variables are reused (if a cycle exists) or how tasks are connected. Another solution is a Jupyter Notebook extension that assigns unique IDs to each cell, making it easier for users to track how cells depend on each other [75]. These tools help manage and simplify the often complex workflows in notebooks.
@@ 5.2.2 Computational Environments in Notebooks
The computational environment refers to the environment in which Jupyter notebooks run [72]. This environment can affect the code execution process because different execution environments may have different dependencies, causing inconsistencies in code behaviour [31]. Researchers have found that data scientists face computational environment-related challenges throughout the analytics workflow, from setting up the notebook to deploying it to production [21].
@@
Although notebooks typically run in a single computational environment, there are advantages to running a notebook in multiple computational environments [25, 31]. For example, parts of a notebook, such as model training algorithms, may require specialized computational resources that are unavailable in a standard environment designed primarily for data exploration or visualization. Cunha et al. [25] presented a solution developed as a Jupyter Notebook extension to automatically determine and migrate selected notebook cells to appropriate computational environments for execution. Their solution leverages abstract syntax trees and dependency tracking to extract the selected notebook cells and the dependencies. Another study discussed the possibility of automating the migration of a computational environment for Jupyter notebooks to a distributed Kubernetes [^12] https://kubernetes.io/ environment [31]. This approach aims to overcome the limitations of Jupyter notebooks, such as scalability and fault tolerance, by dividing notebooks into executable steps and deploying them in a Kubernetes cluster. In addition, ElasticNotebook [88, 86] provides live migration through optimized checkpointing and restoration to reduce migration overhead. Expanding this concept further, Kinanen et al. [70] introduced a custom Jupyter kernel to simplify offloading quantum computation tasks directly from the notebook to remote Kubernetes clusters, thus lowering access barriers to high-performance quantum computing without requiring deep infrastructure expertise.
@@
Recognizing security as a crucial factor in computational environment design, researchers have explored specific strategies for improving Jupyter notebook security. For instance, Lu et al. [92] proposed an architectural security framework leveraging containerization, load balancing, authentication, and encryption mechanisms, thus safeguarding sensitive data within cloud-based notebook execution environments. In addition, Ramsingh and Verma [128] conducted empirical analyses revealing gaps in current notebook security practices, such as poor awareness of threats and insufficient technical expertise. Based on their findings, they proposed the Jupyter multi-layer security (JMLS) defence model explicitly designed to reinforce notebook security. Complementing these perspectives, Cao [16] developed a systematic taxonomy to classify and better understand the variety of network-based security threats impacting Jupyter notebooks, especially in high-performance computing (HPC) scenarios. Their exhaustive analysis highlighted vulnerabilities involving ransomware, data exfiltration, misconfiguration, and resource misuse, providing insights for secure management of computational environments.
@@
Researchers have also investigated various computational environments to tackle Jupyter Notebook reproducibility and usability issues. Sato and Nakamaru [141] addressed reproducibility issues arising from potentially unsafe dynamic notebook modifications by developing Multiverse Notebook, a computational environment that enables safe and efficient “time-travel” (cell-wise checkpointing) based on the POSIX “fork()” mechanism, where a process is assigned to each executed cell and a process tree is maintained as the entire state. To address hidden state and out-of-order execution issues, Weinman et al. [189] created a notebook extension that allows for easy branching (or “forking”) of execution states into separate kernel instances to simplify concurrent exploration of different analytical approaches. Furthermore, Li et al. [87] introduced Kishu, which employs namespace-patching techniques to efficiently revert or undo notebook states without resorting to costly kernel restarts or complete re-executions. To improve Jupyter notebook usability and integration into software development workflows, Titov et al. [160] explored strategies for integrating computational notebooks within IDEs to better align interactive and exploratory notebook usage with standard software engineering practices. Complementarily, acknowledging the cognitive and operational challenges inherent in mixed-methods research workflows, Zhu et al. [204] proposed design concepts for notebook environments that more effectively integrate qualitative and quantitative analysis tasks, thereby streamlining hybrid analytical processes.

### v2

@@ 5.1.3 Reproducibility
Reproducibility ensures that other notebook users can get the same results from a notebook by following the same steps, using the same data, and working in the same environment ahmad2022reproducible. Reproducibility differs from code reuse and cloning, which typically involves adapting code or directly copying code into different environments and applying the code to different data to meet different objectives of users, resulting in outputs that are different from the original notebook.
@@
Reproducibility is known as one of the key promises of Jupyter notebooks, as they are often used to share results with others, allowing users to trace the steps from raw data to final results kluyver2016jupyter. However, achieving reproducibility remains challenging due to issues such as out-of-order execution and repetitive execution of the same cell wang2020assessing. These issues differ from traditional programming, where code execution is typically linear (from top to bottom), ensuring a clear and consistent execution flow. For example, notebook users might execute the same cell multiple times and only save the latest execution state, causing cells located at the notebook’s beginning to be executed after later cells. This nonsequential execution can confuse others trying to reproduce results, affecting notebook reproducibility wang2020assessing, wang2020restoring.
@@
A study by pimentel2019large analyzed more than 1.4 million Jupyter notebooks from GitHub and found significant reproducibility issues. The study noticed that only 25% of the valid notebooks executed without errors under straightforward conditions, and only 5% reproduced results identical to their original outputs. Furthermore, schroder2019reproducible revealed that only 14% of the available notebooks in published academic research on PubMed, [^16] https://pubmed.ncbi.nlm.nih.gov/ where Jupyter notebooks are used as implementation, were reproduced successfully. The study noticed the need for comprehensive documentation and containerization of notebooks, which supports the good practices presented by rule2018ten. In a more recent large-scale analysis, mietchen2025analyzing attempted to automatically rerun thousands of notebooks associated with biomedical research articles in PubMed Central, highlighting many recurring failures (such as missing dependencies and improper execution order) and noting how the corpus of notebooks evolved over time. However, not all notebooks that fail to run are beyond repair. nguyen2025majority revisited prior claims that roughly 76% of public notebooks are non-executable, arguing that this estimate relies on an overly rigid notion of executability. Analyzing 42,546 notebooks, they showed that only about 21% are truly non-restorable. For the remaining notebooks, executability can be improved by 43% through installing missing dependencies and by 28% through generating synthetic data. This suggests that notebooks have higher latent reproducibility than previously reported, provided that tools or users address minor reproducibility pitfalls. We have analyzed the relevant studies and presented a list of good and bad practices that affect the reproducibility of a notebook in Table 6.
@@
Researchers have demonstrated that the reproducibility of Jupyter notebooks can be significantly improved by resolving cell dependencies within the notebook’s code wang2020assessing, wang2021restoring. Cell dependencies represent the relationships between variables, functions, and data in different code cells. wang2020restoring proposed a method to build a cell dependency graph using static analysis to accurately model these relationships and found that 83% of executable notebooks can be reproduced using their approach. samuel2021reproducemegit developed a visualization tool ReproduceMeGit to check the reproducibility of Jupyter notebooks from GitHub by extending the work from pimentel2019large. This tool can automatically install dependencies from requirements.txt or setup.py files and execute the notebook to provide highlights of the reproducibility study through a user interface.
@@ 5.1.4 Provenance
Provenance in notebooks refers to the detailed history of actions, executions, and dependencies that led to the current state of a notebook. It includes the sequence of cell executions, the modifications made to cells, variable dependencies, data flow, and intermediate outputs kerzel2023mlprovlab. Unlike reproducibility which emphasizes achieving identical results under a consistent initial environment, provenance addresses the historical and iterative details present in the notebook’s creation and evolution. Managing provenance in Jupyter Notebook differs from traditional systems due to its iterative and non-linear nature. In a notebook, users can execute cells in any order, creating challenges in tracking the execution history and dependencies [kerzel2021towards]. Empirical observations underscore this challenge, for example, kery2018story found that data scientists often explore ideas non-linearly and later refactor their notebooks into a linear narrative (the “story” in the notebook), which can obscure the trial-and-error process. This highlights the need for provenance tools to capture the actual sequence of analysis steps. Additionally, Jupyter notebooks merge code, data, and outputs into a single file, whereas traditional systems follow a sequential execution process that generates output externally. As traditional systems offer clear execution pipelines and distinct output demonstrations, provenance management becomes more straightforward. In contrast, Jupyter notebooks require specialized tools for managing provenance to capture dynamic state changes and overwritten variables kerzel2023mlprovlab, koop2021notebook.
@@
Researchers developed provenance detection techniques specialized for Jupyter Notebook to reuse notebooks by obtaining the original results. In the early stage of Jupyter Notebook usage, pimentel2015collecting proposed a notebook tool to overcome the limitations in provenance support by automatically capturing and analyzing the execution history and environment settings of code within notebooks. The study extends the noWorkFlow tool murta2015noworkflow developed for Python scripts to capture the provenance of scripts, including control flow information and library dependencies. The proof can be inferred by statically analyzing code cells on execution counts and cell positions within the notebooks koop2021notebook. A scheduler that adjusts the order of cell execution based on data dependencies detected during runtime can be used to improve the accuracy of provenance of Python code in notebooks dao2022runtime.
@@
In addition to focusing on tools and execution kernels, researchers have developed extensions to capture provenance. MLProvLab [kerzel2023mlprovlab] is a JupyterLab extension to track, manage, compare and visualize the provenance of computational experiments. It can capture the provenance at run-time by detecting dependency graphs. Another notebook extension, ProvBook, was designed to capture and visualize the provenance of notebook executions samuel2018provbook. This extension automatically stores and visualizes provenance information within the notebook’s metadata, detailing execution times, inputs, and outputs for each cell. It helps users to see how data and analysis have evolved over time and compare results, which helps to enable notebook reproducibility. kery2018interactions developed an extension named Verdant that records the history of all previous changes in a notebook and quickly retrieves versions of a specific artifact from the existing versions of the entire document. It can help users to visually compare side-by-side multiple versions of different notebook artifacts, including code cells, tables, and images, which are absent in traditional version control systems like Git. Building on the idea of visualizing notebook evolution, eckelt2024loops introduced a system called Loops that leverages provenance to support exploratory data analysis. Loops captures the evolution of a notebook’s state as cells are edited and re-executed, and presents an interactive timeline visualization of these changes. This allows users to replay or inspect how each modification to the code or data affected the results, thereby aiding in understanding the analytical process and preserving intermediate insights.
@@
Expanding on these efforts, researchers have explored versioning mechanisms for notebooks. studtmann2023histree introduced Histree for automatic versioning and visual branching of notebook modifications. It organizes experiment histories into tree-based structures for easier navigation. Similarly, fang2025enhancing proposed a two-dimensional version control approach (realized in a tool called Kishuboard) that captures both code and data state in notebooks. By adjusting “code” and “data” axes (or knobs), users can intuitively branch, roll back, and compare multiple divergent notebook states, achieving flexible exploration without losing past states; user studies showed that such integrated code-and-data versioning significantly enhanced productivity in complex data analysis tasks. Further extending provenance into interactive visual analytics, gadhave2024persist proposed Persist, which captures interaction provenance across code and visualizations. Empirical evaluations showed improvements in analysis efficiency, accuracy, and reproducibility. Finally, Retrograde harrison2024jupyterlab integrates provenance tracking with fairness auditing. It provides real-time, context-aware notifications triggered by notebook events and maintains a dynamic data ancestry graph, helping users detect and address fairness and bias concerns during data preprocessing and modeling.
@@
Summary of code reuse and provenance Code reuse in Jupyter notebooks is essential for improving productivity and collaboration while addressing the challenges associated with their unique structure and iterative workflows. Users can more effectively leverage existing resources and maintain reusable notebooks through techniques such as code cloning, context-aware code search, provenance management, reproducibility practices, and advanced versioning tools. The development of innovative solutions, such as context-sensitive code search tools [li2021nbsearch, li2023edassistant], code dependency resolution systems [wang2020restoring], and provenance tracking extensions [kerzel2023mlprovlab, pimentel2015collecting, harrison2024jupyterlab], highlights ongoing efforts to support effective code reuse and provenance. By integrating these solutions and best practices (see Table 6), notebooks can become more reusable and maintainable, thus establishing a stronger foundation for data science workflows and collaborative research.
@@ 5.2 Managing Computational Environment and Workflow
Managing the computational environment and workflow in Jupyter notebooks is crucial to ensure the successful execution and reproducibility of data science tasks [psallidas2022data, lu2020securing]. Unlike traditional programming environments, which are typically well defined and managed by integrated development environments or build systems, a Jupyter Notebook does not have a separate file to manage the configuration [kluyver2016jupyter]. Traditional programming environments commonly include a configuration file (e.g., “requirements.txt” in Python) to ensure a consistent and reproducible setup for code execution. In contrast, Jupyter notebooks execute code cells independently, and each notebook has its own dependencies [cunha2021context, raghunandan2023measuring]. This section explores various aspects of computational environment management, including empirical studies on notebook workflows, the impact of different environments on code behaviour, and the challenges associated with managing library dependencies.
@@ 5.2.1 Empirical Studies on Workflows
lau2020design analyzed 60 notebooks and summarized four main stages of a data science workflow: importing data into notebooks, writing and editing code, running the code to generate output, and publishing the results. However, workflows can vary. For example, notebooks often start with exploration, where users write lots of code to find interesting patterns in the data, and end with a presentation or explanation [raghunandan2023measuring]. In addition, recent empirical analyses have shed further light on workflow diversity. golendukhina2024unveiling studied 138,376 Kaggle notebooks and identified significant variations in data preprocessing practices correlated with user expertise, highlighting a notable gap between model-centric activities and actual data cleaning efforts. Similarly, zou2024mlpipeline uncovered that data scientists frequently and manually experiment with alternative ML pipeline configurations (e.g., data preparation, model selection), emphasizing the limitations of managing and systematically exploring pipeline variations in traditional notebook environments. On the other hand, psallidas2022data identified two types of pipelines in Jupyter notebooks that are key components of data science workflows: explicit and implicit. Explicit pipelines utilize tools like sklearn.pipeline to define structured steps for tasks such as data pre-processing and model training. In contrast, implicit pipelines rely on ad hoc function calls with libraries such as Pandas to clean, merge, and visualize data [psallidas2022data]. Researchers have also suggested ways to map workflows to better understand how notebooks work. One approach uses directed acyclic graphs to show the flow of data and tasks within a notebook [rehman2019towards]. These graphs can highlight where variables are reused (if a cycle exists) or how tasks are connected. Another solution is a Jupyter Notebook extension that assigns unique IDs to each cell, making it easier for users to track how cells depend on each other [koop2017dataflow]. These tools help manage and simplify the often complex workflows in notebooks. To further understand user behaviour, sparmann2023jugaze introduced an eye-tracking and logging tool named JuGaze which records how users visually navigate notebook cells. This fine-grained analysis of user attention offers insights into workflow patterns and can inform interface designs to better support common usage behaviours.
@@ 5.2.2 Computational Environments in Notebooks
The computational environment in notebooks refers to the environment in which Jupyter notebooks run [kluyver2016jupyter]. This environment can affect the code execution process because different execution environments may have different dependencies, causing inconsistencies in code behaviour [duan2023jup2kub]. Researchers have found that data scientists face computational environment-related challenges throughout the analytics workflow, from setting up the notebook to deploying it to production [chattopadhyay2020whats].
@@
Although notebooks typically run in a single computational environment, there are advantages to running a notebook in multiple computational environments [cunha2021context, duan2023jup2kub]. For example, parts of a notebook, such as model training algorithms, may require specialized computational resources that are unavailable in a standard environment designed primarily for data exploration or visualization. cunha2021context presented a solution developed as a Jupyter Notebook extension to automatically determine and migrate selected notebook cells to appropriate computational environments for execution. Their solution leverages abstract syntax trees and dependency tracking to extract the selected notebook cells and the dependencies. Another study discussed the possibility of automating the migration of a computational environment for Jupyter notebooks to a distributed Kubernetes [^17] https://kubernetes.io/ environment [duan2023jup2kub]. This approach aims to overcome the limitations of Jupyter notebooks, such as scalability and fault tolerance, by dividing notebooks into executable steps and deploying them in a Kubernetes cluster. In addition, ElasticNotebook [li2023elasticnotebook, li2024demonstration] provides live migration through optimized checkpointing and restoration to reduce migration overhead. Expanding this concept further, kinanen2024improving introduced a custom Jupyter kernel to simplify offloading quantum computation tasks directly from the notebook to remote Kubernetes clusters, thus lowering access barriers to high-performance quantum computing without requiring deep infrastructure expertise.
Moreover, bridging notebooks with high-performance computing (HPC) resources has become an area of focus. jayawardana2024enhancing developed a gateway named Cybershuttle that enables seamless integration of personal Jupyter notebooks with remote HPC and cloud infrastructures, allowing researchers to offload heavy computations to supercomputers while using familiar notebook interfaces. In the scientific computing domain, notebooks have been used to write and manage large-scale simulations. For instance, savira2021writing describe strategies for running and analyzing large scientific simulations via Jupyter, and tsai2024libyt introduced Libyt, which integrates the yt analysis library with Jupyter for parallel in situ analysis of simulation data. To improve GPU utilization for interactive model training in notebooks, carver2025notebookos proposed NotebookOS, an operating system for notebooks that replicates kernel processes across multiple GPU servers and dynamically allocates GPUs only during cell execution, significantly reducing idle GPU time and cost in cloud-based notebook services. Additionally, ueno2020ssh developed an SSH-based Jupyter kernel for remote infrastructure administration, enabling users to execute shell commands on remote servers from within a notebook. This extends the notebook’s reach to managing external systems directly through the notebook environment.
@@
Recognizing security as a crucial factor in computational environment design, researchers have explored specific strategies for improving Jupyter notebook security. For instance, lu2020securing proposed an architectural security framework leveraging containerization, load balancing, authentication, and encryption mechanisms, thus safeguarding sensitive data within cloud-based notebook execution environments. In addition, ramsingh2024understanding conducted empirical analyses revealing gaps in current notebook security practices, such as poor awareness of threats and insufficient technical expertise. Based on their findings, they proposed the Jupyter multi-layer security (JMLS) defence model explicitly designed to reinforce notebook security. Complementing these perspectives, cao2024jupyter developed a systematic taxonomy to classify and better understand the variety of network-based security threats impacting Jupyter notebooks, especially in high-performance computing (HPC) scenarios. Their exhaustive analysis highlighted vulnerabilities involving ransomware, data exfiltration, misconfiguration, and resource misuse, providing insights for secure management of computational environments.
@@
Researchers have also investigated various computational environments and techniques to tackle Jupyter Notebook reproducibility and usability issues. sato2024multiverse addressed reproducibility issues arising from potentially unsafe dynamic notebook modifications by developing Multiverse Notebook, an environment (initially introduced as a prototype by christophe2023moon that enables safe and efficient “time-travel” (cell-wise checkpointing) based on the POSIX “fork()” mechanism. In this design, each executed cell runs as a separate process and the collection of processes forms a tree representing the entire notebook state. To address hidden state and out-of-order execution issues, weinman2021fork created a notebook extension that allows for easy branching (or “forking”) of execution states into separate kernel instances to simplify concurrent exploration of different analytical approaches. Furthermore, li2024kishu introduced Kishu, which employs namespace-patching techniques to efficiently revert or undo notebook states without resorting to costly kernel restarts or complete re-executions. Complementing these approaches, zhang2019juneau presented Juneau, which repurposes the notebook backend as a data lake management layer that integrating provenance-aware indexing and reuse of code and data artifacts directly into Jupyter to enhance usability.
To improve Jupyter notebook usability and integration into software development workflows, titov2024hidden explored strategies for integrating computational notebooks within IDEs to better align interactive and exploratory notebook usage with standard software engineering practices. Similarly, kim2024japper presented Japper, a framework that streamlines the conversion of Jupyter notebooks into standalone scientific web applications, thereby simplifying the transition from ad hoc analyses to deployable, interactive tools. In addition, acknowledging the cognitive and operational challenges inherent in mixed-methods research workflows, zhu2024facilitating proposed design concepts for notebook environments that more effectively integrate qualitative and quantitative analysis tasks, thereby streamlining hybrid analytical processes.